# Understanding and Mitigating the Impacts of Differentially Private Census Data on State Level Redistricting

### v1

@@ Abstract
Data from the Decennial Census is published only after applying a disclosure avoidance system (DAS). Data users were shaken by the adoption of differential privacy in the 2020 DAS, a radical departure from past methods. The change raises the question of whether redistricting law permits, forbids, or requires taking account of the effect of disclosure avoidance. Such uncertainty creates legal risks for redistricters, as Alabama argued in a lawsuit seeking to prevent the 2020 DAS’s deployment. We consider two redistricting settings in which a data user might be concerned about the impacts of privacy preserving noise: drawing equal population districts and litigating voting rights cases. What discrepancies arise if the user does nothing to account for disclosure avoidance? How might the user adapt her analyses to mitigate those discrepancies? We study these questions by comparing the official 2010 Redistricting Data to the 2010 Demonstration Data—created using the 2020 DAS—in an analysis of millions of algorithmically generated state legislative redistricting plans. In both settings, we observe that an analyst may come to incorrect conclusions if they do not account for noise. With minor adaptations, though, the underlying policy goals remain achievable: tweaking selection criteria enables a redistricter to draw balanced plans, and illustrative plans can still be used as evidence of the maximum number of majority-minority districts that are possible in a geography. At least for state legislatures, Alabama’s claim that differential privacy “inhibits a State’s right to draw fair lines” appears unfounded.
@@ 1 Introduction
Concerns persisted even after the data were published. In early 2022, one of us was contacted by lawyers at the ACLU in connection with racial gerrymandering cases in Georgia and Arkansas. They wondered whether the new disclosure avoidance system could affect the reported number of majority-Black districts in a demonstration redistricting plan. [^1] Ultimately, this issue was not raised by defendants. No expert testimony was needed. If so, the States could try to take advantage of this uncertainty at trial.
@@
In contrast, the 2020 DAS is based on differential privacy, a modern framework for formally quantifying and limiting such disclosure risks [4, 5]. The 2020 DAS synthesizes and aggregates a new set of microdata census responses that have no one-to-one correspondence with the actual microdata in the CEF, but that closely approximate many of its statistics. While previous DASes have also introduced error into census tabulations, the 2020 DAS is the first to perturb total population counts in every geography smaller than states. This is inherent for the formal guarantees sought by the Census Bureau. Relative to the population, these errors can be quite large in very small geographies (24.79% in urban census blocks) but much smaller even in moderately-sized geographies (0.31% for counties of less than 1,000 people).
@@
Many viewed the 2020 DAS as undermining the legal and political legitimacy of the the data’s use for redistricting [6, 7, 8], calling for a rethinking of disclosure avoidance in the census. Others disagreed [9, 10]. The argument against the new DAS was made most forcefully by Kenny et al. [6]. They analyze the impact of the 2020 DAS using an ensemble-based redistricting analysis: a modern, computationally-intensive approach to exploring properties of the space of reasonable redistricting plans in a given geography. They conclude that “nonrandom local errors can aggregate into substantively large and unpredictable biases at district levels” and claim “the added noise makes it impossible to follow the principle of One Person, One Vote, as it is currently interpreted by courts and policy-makers” [6].
@@
The concern is easy to understand: two districts whose populations are equal according to the official 2020 Redistricting Data will very likely have different populations according to the confidential 2020 CEF. If such discrepancies are large or biased, they could impact electoral power. And if the Supreme Court’s One-Person, One-Vote (OPOV) caselaw requires balancing populations according to the CEF—as Alabama argued in its lawsuit—then redistricting plans created using the Redistricting Data could be illegal.
At the heart of this paper is a legal question. What data should redistricting law treat as the ground truth: the data as enumerated (the CEF) or as published (the Redistricting Data)? No court has ever faced this question, and we do not know how a court might decide. Alabama’s lawsuit assumes that the enumerated data are treated as ground truth. Kenny et al. [6] consider this view to be the law as “currently interpreted by courts and policy-makers.” Cohen et al. [9] argue the opposite: that the official Redistricting Data should be used as the ground truth. They see value in maintaining the “legal fiction” of perfect accuracy when the noise from disclosure avoidance is small relative to other data quality issues (quoting Georgia v. Ashcroft (2003)).
We won’t answer the question. Instead, we ask how treating the enumerated data as the ground truth would affect a redistricter’s ability to comply with the law. Our results may help redistricters better understand and mitigate the litigation risk arising from the legal uncertainty.
@@
This paper considers a data user who only has access to the officially published 2020 Redistricting Data (after disclosure avoidance), but is concerned that the new DAS will substantially alter their ability to comply with redistricting requirements as measured using the confidential CEF data (before disclosure avoidance). [^2] As explained in Section 4, we do not have access to CEF and must use a different dataset as a stand in. See Section 7.1 for a discussion of how this does or does not affect our results. We ask whether the user is able to achieve two goals. First, to produce plans balanced within One Person, One Vote population tolerances. Second, to determine how many majority-minority districts (MMDs)—where at least 50% of the population belongs to the relevant minority group—are possible in a given geography. We study both questions at the state-legislative level, where the legal standing of such plans depend on sharp thresholds (see Section 2). For OPOV, population deviations less than 10% are presumptively constitutional; larger deviations are presumptively unconstitutional. And showing that a political geography admits some number of majority-minority districts—districts with at least 50% of the population in the relevant minority group—is necessary to successfully overturn racially gerrymandered plans.
Main findings
Overall, our results suggest that one can use the 2020 Redistricting Data to meet two policy-relevant threshold requirements (OPOV, MMDs) on the unseen data. While discrepancies do arise, they often take a predictable form and are straightforward to account for. We support our claims with a large-scale ensemble analysis (Section 4) and discuss important limitations of our analysis in Section 7.1.
In the context of One-Person One-Vote, we observe large discrepancy rates in the fraction of plans satisfying a strict 5% population deviation limit3335% deviation as measured using the method of [6] and others, which approximates the 10% deviation threshold as measured by courts (see Section 4.1). across many state legislative geographies (greatly extending a finding of [6]). Among plans sampled with at most 5% population deviation on the noised data, the median (across legislative geographies) discrepancy rate—the fraction of plans exceeding 5% deviation on the unseen data—was 9%. The max was over 80%. However, using a population deviation threshold in the redistricting sampling algorithm that is slightly tighter than the 5% policy target is an effective countermeasure. Sampling with a 4.8% limit makes the same discrepancy rate fall below 2% for 75% of geographies examined. Sampling with a 4% limit does so for all geographies examined. These results are presented in Section 5.1.
We focus our examination of MMDs on majority Black districts. We observe that discrepancy in the apparent number of majority Black districts in a plan can be large and biased. On average, the plans that result from optimizing the number of majority Black districts in the published data have fewer majority Black districts in the unseen data. Moreover, the magnitude of the discrepancy varies greatly by geography. Nevertheless, we find the published data still provide reliable evidence of the maximum number of majority Black districts achievable within a given state. This is the quantity most relevant to the Gingles 1 test in racial gerrymandering cases. These results are presented in Section 6.1.
Along the way, we develop a richer qualitative understanding of the ways that noise from disclosure avoidance affects these policy-relevant quantities. We put these effects in context, showing that in the settings we consider the error introduced by the 2020 DAS is small relative to both total population and other sources of noise.
@@ 2.1 One Person, One Vote
The One-Person, One-Vote principle holds that any two people living in the same state should have about the same voting power. In Karcher v. Daggett, the Supreme Court held that “congressional districts [must] be apportioned to achieve population equality as nearly as is practicable.” Moreover, the decennial census “furnishes the only basis” for balancing populations, despite it’s imperfections. In practice, this often leads to Congressional districts being balanced as closely as possible according to decennial census releases, despite the Court acknowledging the “inherent artificiality” of the data. After the 2010 Census, 26 states balanced their congressional districts to within a single person in total population, although larger deviations have been allowed [11]. State-level legislative districts must also be balanced in total population, although the requirement is much less strict. For state legislative districts, population deviations below 10% “will ordinarily be considered de minimis,” while “disparities larger than 10% creates a prima facie case of discrimination” (Brown v. Thomson 1983). In practice, state legislative population deviations are often much smaller than 10%. One reason we do not study Congressional redistricting is that there is no amount of population deviation that is considered de minimis (Karcher v. Daggett). As such, there is no policy-relevant numerical target that a redistricter can hope to achieve after accounting for noise.
Note that courts define population deviation as the relative difference between the largest and smallest districts. We instead adopt the convention of [6] and others, using the maximum relative difference from the ideal district population. A 10% deviation under the former definition is approximately equal to a 5% deviation under the definition used in this paper (see Sec. 4.4.1).
@@ 2.2 Voting Rights Act
Section 2 of the Voting Rights Act of 1965 outlaws vote dilution on the basis of race, including racial gerrymandering. The Supreme Court laid out a framework for judging racial gerrymandering claims in Thornburg v. Gingles (1986), and recently reaffirmed the framework in Allen v. Milligan (2023). To successfully challenge an enacted map, a minority group must first pass a three-part threshold test: the Gingles preconditions. The second and third preconditions together require voting to be polarized to such an extent that the minority group is prevented from electing its candidates of choice (Thornburg v. Gingles 1986).
Our paper focuses specifically on the first Gingles precondition, Gingles 1. Gingles 1 requires the minority group to be “sufficiently large and geographically compact to constitute a majority in a reasonably configured district” (Allen v. Milligan 2023). This is typically accomplished by producing for the court one or more demonstration maps (eleven in Allen), each with more majority-minority districts than the enacted map being challenged. Gingles 1 is satisfied if at least one demonstration map is reasonably configured (“if it comports with traditional districting criteria”). Whether a district is majority-minority is based on the voting-age population rather than the total population (Bartlett v. Strickland, 2009).
Throughout this paper, we specifically focus on Black voters. This reflects a common scenario in voting rights litigation and the largest racial minority in the USA according to Census data. Hence, districts will be considered majority-minority if a majority of the voting age population is Black (including respondents who selected multiple races).
@@ 2.3 Redistricting with imperfect data
The question of which dataset is the proper basis for redistricting law is one facet how redistricting law deals with uncertainty or error more generally [12]. It has long been recognized that the Decennial Census is far from perfect. For example, significant overage error that varies race: the 2020 Census was measured to have a 3.30% Black undercount, a 4.99% Hispanic undercount, and a 0.66% White overcount [13]. Even if the Decennial Census was perfect, populations of areas can change significantly in the 10 years between censuses when districts are largely unchanged.
@@
The Supreme Court recognizes these issues. Generally, the Redistricting Data is considered “the only basis for good-faith attempts to achieve population equality” despite being “less than perfect” (Karcher v. Daggett 1983). But the Court has also acknowledged that deviating from the Redistricting Data is sometimes appropriate. In Mahan v. Howell (1973), a plan was considered malapportioned when about 18,000 people were known to live off of the naval base in which they were counted in the 1970 Census.
3 Data and methods
3.1 Decennial Census and disclosure avoidance
@@
We use data from the 2010 Census P.L. 94-171 Redistricting Data Summary Files (henceforth SWAP) and the Privacy-Protected 2010 Census Demonstration Data Vintage 2021-06-08 (DEMO) [14]. SWAP is the official redistricting data from the 2010 Census, produced by applying the swapping-based 2010 DAS to the 2010 CEF. DEMO was produced by applying the 2020 DAS to the same 2010 CEF. These data files contain tabulations of population by voting age, sex, race, ethnicity for each geographic unit on the census geographic hierarchy—census blocks, block groups, tracts, counties, states, and the nation as a whole. SWAP and CEF agree on the total and voting age populations of every census geography, as the 2010 DAS held these statistics as invariant, but other demographic counts may differ. In contrast, DEMO and CEF agree on the total population only at the state level, and all sub-state counts may differ. We link these data with TIGER/Line Shapefiles [15] and block equivalency files for legislative districts enacted after the 2010 Census [16] provided by the Census Bureau to create the input data for our MCMC redistricting algorithms.

### v2

@@ Abstract
Data from the Decennial Census is published only after applying a disclosure avoidance system (DAS). Data users were shaken by the adoption of differential privacy in the 2020 DAS, a radical departure from past methods. The goal of this paper is to better understand how the perturbations from the 2020 DAS combine with sharp legal thresholds to impact redistricting. We consider two redistricting settings in which a data user might be concerned about the impacts of privacy preserving noise: drawing equal population districts and litigating voting rights cases. What discrepancies arise if the user does nothing to account for disclosure avoidance? How can the discrepancies be understood and accounted for? We study these questions by comparing the official 2010 Redistricting Data to the 2010 Demonstration Data—created using the 2020 DAS—in an analysis of millions of algorithmically generated state legislative redistricting plans. We find that thresholding can amplify the impact of the noise from disclosure avoidance. Large discrepancies do occur, but in ways that are well-captured by simple models and appear to be possible to account for. We demonstrate the utility of these models by proposing an approach to mitigate discrepancies when balancing district populations. At least for state legislatures, Alabama’s claim that differential privacy “inhibits a State’s right to draw fair lines” lacks support.
@@ 1 Introduction
Concerns persisted even after the data were published. In early 2022, one of us was contacted by lawyers at the ACLU in connection with racial gerrymandering cases in Georgia and Arkansas. They wondered whether the new disclosure avoidance system could affect the reported number of majority-Black districts in a demonstration redistricting plan. [^1] Ultimately, this issue was not raised by defendants. No expert testimony was needed.
@@
In contrast, the 2020 DAS is based on differential privacy, a modern framework for formally quantifying and limiting such disclosure risks [17, 1]. The 2020 DAS synthesizes and aggregates a new set of microdata census responses that have no one-to-one correspondence with the actual microdata in the CEF but closely approximate many of its statistics. While previous DASes have also introduced error into census tabulations, the 2020 DAS is the first to perturb total population counts in every geography smaller than states. This is necessary for the formal guarantees sought by the Census Bureau. Relative to the population, these errors can be quite large in very small geographies (24.79% in urban census blocks) but are much smaller even in moderately-sized geographies (0.31% for counties of less than 1,000 people).
@@
Many viewed the 2020 DAS as undermining the legitimacy of the data’s use for redistricting and called for a rethinking of disclosure avoidance in the census [24, 37, 47, 5]. Others disagreed [12, 23]. The argument against the new DAS was made most forcefully by Kenny et al. [24]. They analyze the impact of the 2020 DAS using an ensemble-based redistricting analysis: a modern, computationally-intensive approach to exploring properties of the space of reasonable redistricting plans in a given geography. They conclude that “nonrandom local errors can aggregate into substantively large and unpredictable biases at district levels” and claim “the added noise makes it impossible to follow the principle of One Person, One Vote, as it is currently interpreted by courts and policy-makers” [24].
@@
The concern is easy to understand: two districts whose populations are equal according to the official 2020 Redistricting Data will very likely have different populations according to the confidential 2020 CEF. If such discrepancies are large or biased, they could impact electoral power.
Sharp legal thresholds can amplify discrepancies, turning a continuous source of error into a discrete category change. For example, 10% population deviation is the cutoff between a state legislative plan being presumed discriminatory or not. Prior work found that about 5% of plans generated for the Louisiana state senate using the 2020 Redistricting Data would have exceeded that population deviation threshold according to the 2020 CEF [24].
This is what Alabama and Kenny et al. [24] are worried about: that plans created using the Redistricting Data could turn out to be illegal—potentially “impeding” or even “inhibiting” the ability to “draw fair lines,” as Alabama put it. (We question the underlying legal theory, as discussed in Section 8.) Despite prior work quantifying the magnitude and characteristics of the perturbations, little is known about the interaction with threshold redistricting requirements.
@@
The goal of this paper is to better understand how perturbations from disclosure avoidance combine with sharp legal thresholds to impact redistricting. What sorts of discrepancies arise and why? Can they be mitigated?
To that end, we consider a conscientious redistricter who wants to minimize the impact of disclosure avoidance on their map, even if not legally required. Our redistricter only has access to the officially published 2020 Redistricting Data (after disclosure avoidance), but wants their maps to comply with threshold redistricting requirements as measured using the confidential CEF data (before disclosure avoidance). [^2] As explained in Section 3.1, we do not have access to CEF and must use a different dataset SWAP as a stand in. See Section 7 for a discussion of how this does and does not affect our results.
We ask whether there exists methods by which the redistricter is able to achieve two goals, measured against the confidential CEF. First, to produce plans balanced within One Person, One Vote (OPOV) population tolerances. Second, to produce plans with as many majority-minority districts (MMDs) as possible.
@@
We study both questions at the state legislative level, where the legal standing of such plans depend on sharp thresholds (see Section 2). For OPOV, population deviations less than 10% are presumptively constitutional; larger deviations are presumptively unconstitutional. And showing that a political geography admits some number of majority-minority districts—districts with at least 50% of the population in the relevant minority group—is necessary to successfully overturn racially gerrymandered plans. We focus our examination of MMDs on majority-Black districts, reflecting a common scenario in voting rights litigation and the largest racial minority in the USA.
Overview of Methods
This paper employs ensemble analysis. We algorithmically sample large collections of possible districting plans and observe how statistics of our ensembles differ between our two data sources (see Section 3 for details).
To study the effect of the DAS, one would ideally compare the data with and without disclosure avoidance, namely the 2020 CEF and 2020 Redistricting Data. However, the CEF is confidential to those outside of the Census Bureau. We instead use two Census datasets which we call SWAP and DEMO as stand-ins for the 2020 CEF and Redistricting Data, respectively. We are using these datasets exactly as intended: The Census Bureau created DEMO to help stakeholders study the impacts of the 2020 DAS by comparing it to SWAP. See Section 3 for more information and Section 7 for further discussion on the limitations of using these datasets.
Our high-level approach is simple, and is adapted from Kenny et al. [24]. First, we generate an ensemble of many redistricting plans. These plans are drawn to satisfy some constraint—or to maximize some objective—on the DEMO dataset (standing in for 2020 Redistricting Data). Second, across plans in the ensemble, we see how often the constraint is violated—or how the value of the objective compares—on the SWAP dataset (standing in for 2020 CEF).
Contributions and Main Findings
We investigate the discrepancies that occur when maps created using data perturbed by the 2020 disclosure avoidance system are measured against sharp threshold requirements using the unseen, unperturbed dataset. Overall, we find that thresholding can amplify the impact of disclosure avoidance noise, in ways that are well-captured by simple models and possible to account for. In particular, there exists a method to satisfy population balance requirements on SWAP using only the DEMO data with confidence. We also provide simple probabilistic models that help explain our empirical observations.
Below is a summary of our contributions and some of our main findings. Our quantitative findings are specific to our experimental design (e.g., state legislative districts, using precincts as building blocks); drawing quantitative conclusions for other redistricting settings (e.g., smaller sub-state districts, incorporating different constraints or objectives) would require adapting our techniques. We discuss these and other important limitations in Section 7.
For One Person, One Vote (Section 5)
We extend the ensemble analysis in prior work [24] from one state legislative geography to 93. As in the prior work, we observe significant discrepancies. For example, among plans sampled with at most 5% population deviation on DEMO, the discrepancy rate—the fraction of plans exceeding 5% deviation on SWAP—was at least 40% in half of the legislative geographies we examined.3335% deviation as measured using the method of [24] and others, which approximates the 10% deviation threshold as measured by courts (see Section 3.3).
At the same time, there exists a method to draw districts that can substantially reduce or eliminate observed threshold violations within our district-generation procedure. The countermeasure is simple: use a population deviation threshold in the redistricting sampling algorithm that is slightly tighter than the 5% policy target (i.e., applying an offset). Sampling with a 4.5% limit yields a discrepancy rate of 0% for over half of geographies examined. Sampling with a 4% limit does so for 90% geographies examined.
For majority-minority districts (Section 6)
We extend the ensemble analysis in prior work [24] from one state legislative geography to 52. We observe significant discrepancies, but in the opposite direction than prior work (see Section 4). For example, among Georgia state house plans sampled using DEMO, 9% have a different number of majority-Black districts on SWAP—almost always fewer.
When sampling plans using a method that maximizes the number of majority-Black districts (as is common in Voting Rights Act litigation), this jumps to 66%.
Simple models for understanding discrepancies
In both cases, we give simple probabilistic models that help explain many of the observed phenomena, including the effectiveness of the mitigation proposed for achieving population balance and the impact of maximizing the number of majority-Black districts on the discrepancies that result. Our hope is that these models prove useful for thinking about the impact of disclosure avoidance noise in settings beyond our specific experimental design.
@@ 2.1 One Person, One Vote
The One-Person, One-Vote principle holds that any two people living in the same state should have about the same voting power. This means that the total population of different districts must be balanced according to the decennial census data (despite its imperfections, see Sec. 2.3).
For Congressional districts, “districts [must] be apportioned to achieve population equality as nearly as is practicable” (Karcher v. Daggett). No amount of population deviation is considered de minimis. In practice, many states balance congressional districts to within a single person in total population. This is one reason that Congressional redistricting is not the focus of our study: there is no policy-relevant numerical target that a redistricter can hope to achieve after accounting for noise.
State-level legislative districts must also be balanced in total population, although the requirement is much less strict. For state legislative districts, population deviations below 10% “will ordinarily be considered de minimis,” while “disparities larger than 10% creates a prima facie case of discrimination” (Brown v. Thomson 1983). In practice, state legislative population deviations are often much smaller than 10%.
Note that courts define population deviation as the relative difference between the largest and smallest districts. We instead adopt the convention of [24] and others, using the maximum relative difference from the ideal district population. A 10% deviation under the former definition is approximately equal to a 5% deviation under the definition used in this paper (see Sec. 3.3).
@@ 2.2 Voting Rights Act
Section 2 of the Voting Rights Act of 1965 outlaws vote dilution on the basis of race, including racial gerrymandering. The Supreme Court laid out a framework for judging racial gerrymandering claims in Thornburg v. Gingles (1986), and recently reaffirmed the framework in Allen v. Milligan (2023).
To successfully challenge an enacted map, a minority group must first pass a three-part threshold test: the Gingles preconditions. Our paper focuses on the first Gingles precondition, Gingles 1. (The second and third preconditions together require voting to be polarized to such an extent that the minority group is prevented from electing its candidates of choice.)
Gingles 1 requires the minority group to be “sufficiently large and geographically compact to constitute a majority in a reasonably configured district” (Allen v. Milligan, 2023). This is typically accomplished by producing for the court one or more demonstration maps (eleven in Allen), each with more majority-minority districts than the enacted map being challenged. Gingles 1 is satisfied if at least one demonstration map is reasonably configured (“if it comports with traditional districting criteria”).
Throughout this paper, we focus on Black voters. This reflects a common scenario in voting rights litigation and the largest racial minority in the USA according to Census data. Hence, districts will be considered majority-minority if a majority of the voting age population is Black (Bartlett v. Strickland, 2009), excluding those who selected multiple races.
Gingles claims are also used to allege vote dilution against other minority groups. We present corresponding results for majority-Hispanic districts in Table 7 in Appendix A.
@@ 2.3 Redistricting with Imperfect Data
It has long been recognized that the Decennial Census is far from perfect [49]. For example, coverage errors are large and vary by race: estimates for 2020 are 3.30% Black undercount and 4.99% Hispanic undercount, compared to 0.66% White overcount [28]. Even if the Decennial Census was a perfect record of the population on Census Day, populations can change significantly in the 10 years between censuses when districts have historically remained unchanged.
@@
The Supreme Court recognizes these issues. Generally, the Census’s official Redistricting Data is considered “the only basis for good-faith attempts to achieve population equality” despite being “less than perfect” (Karcher v. Daggett 1983). But the Court has also acknowledged that deviating from the Redistricting Data can be appropriate. In Mahan v. Howell (1973), a plan was considered malapportioned when about 18,000 people were known to live off of the naval base in which they were counted in the 1970 Census.
Alabama’s lawsuit against the 2020 DAS raises an important question. What data should redistricting law treat as the ground truth: the data as collected, edited, and imputed (the CEF) or as published (the Redistricting Data)? No court has ever faced this precise question.
Kenny et al. [24] consider Alabama’s view—treat CEF as ground truth for redistricting—to be the law as “currently interpreted by courts and policy-makers.” Under this view, discrepancies between the Redistricting Data and the CEF could plausibly impede redistricting by forcing mapmakers to account for measurement error, and might even be argued to inhibit it if no plan could satisfy threshold requirements when judged against the CEF. (Looking ahead, our results suggest that for the state-legislative setting and plan-generation procedures we study, the “impede” concern is real but manageable; see Section 8 for further discussion).
We agree instead with Cohen et al. [12]: the Redistricting Data should remain the only benchmark for redistricting requirements, considering the overall scale of the noise from disclosure avoidance in the face of other sources of imperfection that the Supreme Court has long countenanced, and the infeasibility of enforcing requirements measured against a confidential data source. Under this view, redistricters would not be required to account for data discrepancies to meet legal threshold requirements, though they may still wish to do so as a prudential matter.
3 Methods and Data
3.1 Data Sources
We use data from the 2010 Census P.L. 94-171 Redistricting Data Summary Files (henceforth SWAP) and the Privacy-Protected 2010 Census Demonstration Data Vintage 2021-06-08 (DEMO) [45].
@@ 3.1 Data Sources
SWAP is the official redistricting data from the 2010 Census, produced by applying the swapping-based 2010 DAS to the 2010 CEF. DEMO was produced by applying the 2020 DAS to the same 2010 CEF. Both DEMO and SWAP contain tabulations of population by voting age, sex, race, ethnicity for each geographic unit on the census geographic hierarchy—census blocks, block groups, tracts, counties, states, and the nation as a whole.
SWAP and CEF agree on the total and voting age populations of every census geography, and agree on all statistics at the state level. In contrast, DEMO and CEF agree on the total population only at the state level, and all sub-state counts may differ. Neither SWAP nor DEMO agree with the CEF on population by race for census blocks.
To create the input data for our redistricting algorithms, we link these data with TIGER/Line Shapefiles [43] and block equivalency files for voting precincts and legislative districts enacted after the 2010 Census [42]. For precincts, we additionally joined our data with dual graphs provided by the maintainers of the GerryChain software package [36] for simpler integration into our algorithm.